# A Deep Learning Method for Simultaneous Denoising and Missing Wedge Reconstruction in Cryogenic Electron Tomography

### v2

@@ 2.1.2 Related methods
As mentioned before, DeepDeWedge is closely related to noise2noise-based denoising methods and the denoising and missing wedge-filling method IsoNet [Liu+22].
@@
Noise2noise-like methods for cryo-ET as implemented in CryoCARE [Buc+19] or Warp [TC19] take one or more tilt series as input and return denoised tomograms. A randomly initialized network $f_{\text{$\boldsymbol{\theta}$}}$ is fitted for denoising on sub-tomograms of FBP reconstructions $\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{t}^{\text{0}}})$ and $\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{t}^{\text{1}}})$ of sub-tilt-series ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{t}^{\text{0}}}$ and ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{t}^{\text{1}}}$ obtained from a full tilt series $\mathbf{t}$. The model is fitted by minimizing a loss function (typically the mean-squared error), between the output of the model $f_{\text{$\boldsymbol{\theta}$}}$ applied to one noisy sub-tomogram and the corresponding other noisy sub-tomogram. The fitted model is then used to denoise the two reconstructions $\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{t}^{\text{0}}})$ and $\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{t}^{\text{1}}})$, which are then averaged to obtain the final denoised tomogram.
Contrary to the denoising methods above, DeepDeWedge fits a network that is incentivized to not only denoise but also to fill the missing wedge.
@@
IsoNet takes a small set of, say, one to ten tomograms, and produces denoised, missing-wedge corrected versions of those tomograms. Similar as for DeepDeWedge, a randomly initialized network is fitted for denoising and missing wedge reconstruction on sub-tomograms of these tomograms, however the fitting process is different. Inspired by noisier2noise, the model is fitted on the task of mapping sub-tomograms that are further corrupted with an additional missing wedge and additional noise onto their non-corrupted versions. After each iteration, the intermediate model is used to predict the content of the original missing wedges of all sub-tomograms. The predicted missing wedge content is inserted into all sub-tomograms, which serve as input to the next iteration of the algorithm.
@@
On the contrary, our denoising approach is noise2noise-like, as in CryoCARE. This leads to better performance, as we will see later, as well as requiring fewer assumptions and no hyperparameter tuning. Specifically, noisier2noise-like denoising requires knowledge of the noise model and strength (see Appendix A.2). As this knowledge is typically unavailable, [Liu+22, ] propose approximate noise models from which the user has to choose. After model fitting, the user must manually decide which iteration and noise level gave the best reconstruction. Thus, IsoNet’s noisier2noise-inspired denoising approach requires several hyperparameters for which good values exist but are unknown. Therefore, IsoNet requires tuning to achieve good results. Our denoising approach introduces no additional hyperparameters and does not require knowledge of the noise model and strength.
@@
The main commonality between DeepDeWedge and IsoNet is the noisier2noise-like mechanism for missing wedge reconstruction, which consists of artificially removing another wedge from the sub-tomograms and fitting the model to reconstruct the wedge. Moreover, like IsoNet, DeepDeWedge fills in the missing wedges of the model inputs. In IsoNet, one also has to fill in the missing wedges of the model targets. This is necessary because, contrary to our loss $\ell$ form Equation (2), IsoNet’s loss function does not ignore the targets’ missing wedges via masking in the Fourier domain.
2.1.3 Theoretical Motivation
@@
Figure 2: Illustration of a single ground-truth structure $\mathbf{v}^{*}$, and three model input-target pairs corresponding to three different random positions of the original missing wedge (blue) for the setup of our theoretical motivation. We visualize a 2D structure with no additive noise and random in-plane rotations for simplicity. The small inset boxes in each patch show the absolute values of the Fourier transforms of the images and regions that are zeroed out by the missing wedge masks.
@@
Here, we present a theoretical result that motivates the choice of our per-sample loss $\ell$ defined in Equation (2). The discussion in this section does not consider the heuristic of updating the missing wedges of the model inputs which is part of DeepDeWedge’s model fitting described in Section 2.1.1. Moreover, we consider an idealized setup that deviates from practice to motivate our loss.
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}}=\mathbf{F}^{-1}\mathbf{M}\mathbf{F}\big{(}\mathbf{v}^{*}+\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{n}^{\text{0}}}}\big{)},\qquad{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{1}}}=\mathbf{F}^{-1}\mathbf{M}\mathbf{F}\big{(}\mathbf{v}^{*}+\mathbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{n}^{\text{1}}}}\big{)},$ (4)
@@
where $\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{n}^{\text{0}}}},\mathbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{n}^{\text{1}}}}\in\mathbb{R}^{N\times N\times N}$ are random noise terms and $\mathbf{M}\in\{0,1\}^{N\times N\times N}$ is the missing wedge mask. From the first measurement, we generate a noisier observation ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}}=\mathbf{F}^{-1}\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}\mathbf{F}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}}$ by applying a second missing wedge mask $\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}$. The noisier observation has two missing wedges: the wedge introduced by the first and the wedge introduced by the second mask. We assume that the two masks follow a joint and symmetric distribution. An example of a choice of such distributions is that for each mask, a random wedge is chosen uniformly at random.
Figure 2 illustrates three datapoints in 2D. The bottom row right next to the ground truth shows missing wedge-corrupted versions of the ground truth. For clarity, we set the additive noise to zero, which implies ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{1}}}$. The top row shows noisier observations, i.e. observations with two missing wedges, of the samples in the second row, which have only one missing wedge.
@@
We then train a neural network $f_{\text{$\boldsymbol{\theta}$}}$ to minimize the loss
@@
$\mathcal{L}(\text{$\boldsymbol{\theta}$})=\mathop{\mathbb{E}}_{\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{n}^{\text{0}}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{n}^{\text{1}}}}\left[\left|\left|\big{(}\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}\mathbf{M}+2\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}^{C}\mathbf{M}\big{)}\mathbf{F}\Big{(}f_{\text{$\boldsymbol{\theta}$}}\big{(}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}}\big{)}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{1}}}\Big{)}\right|\right|_{2}^{2}\right],$ (5)
@@
After training, we can use the network to estimate the ground-truth volume by applying the network to another noisy observation ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}}$. The following proposition establishes that this is equivalent to training the network on a supervised loss to reconstruct the input ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}}$, provided the two masks are non-overlapping.
@@
Assume that the noise $\mathbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{n}^{\text{1}}}}$ is zero-mean and independent of the noise $\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{n}^{\text{0}}}}$, and of the masks $(\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}})$, and assume that the noise $\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{n}^{\text{0}}}}$ is also independent of the masks. Moreover, assume that the joint probability distribution $P$ of the missing wedge masks $\mathbf{M}$ and $\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}$ is symmetric, i.e., $P(\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}})=P(\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}},\mathbf{M})$, and that the missing wedges do not overlap. Then the loss $\mathcal{L}$ is proportional to the supervised loss
$\mathcal{R}(\text{$\boldsymbol{\theta}$})=\mathop{\mathbb{E}}_{\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{n}^{\text{0}}}}\left[\left|\left|f_{\text{$\boldsymbol{\theta}$}}\big{(}\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}}}\big{)}-\mathbf{\mathbf{v}_{missing}^{*}}\right|\right|_{2}^{2}\right],$ (6)
i.e., $\mathcal{L}(\text{$\boldsymbol{\theta}$})=\mathcal{R}(\text{$\boldsymbol{\theta}$})+c,$ where $c$ is a numerical constant independent of the network parameters $\boldsymbol{\theta}$.
The proof of Proposition 1 is in Appendix C, where we consider the more general case where the masks $\mathbf{M}$ and $\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}$ may overlap, resulting in a slightly different supervised loss $\mathcal{R}$.

### v3

@@ 2.2 Related Work
DeepDeWedge builds on Noise2Noise-based denoising methods and is related to the denoising and missing wedge-filling method IsoNet [4]. We first discuss the relation between DeepDeWedge and Noise2Noise-based methods, which do not reconstruct the missing wedge.
@@
Noise2Noise-based denoising algorithms for cryo-ET as implemented in CryoCARE [11] or Warp [13] take one or more tilt series as input and return denoised tomograms. A randomly initialized network $\mathrm{f}_{\text{$\bm{\theta}$}}$ is fitted for denoising on sub-tomograms of FBP reconstructions $\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{t}^{\text{0}}})$ and $\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{t}^{\text{1}}})$ of sub-tilt-series ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{t}^{\text{0}}}$ and ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{t}^{\text{1}}}$ obtained from a full tilt series $\mathbf{t}$. The model is fitted by minimizing a loss function (typically the mean-squared error), between the output of the model $\mathrm{f}_{\text{$\bm{\theta}$}}$ applied to one noisy sub-tomogram and the corresponding other noisy sub-tomogram. The fitted model is then used to denoise the two reconstructions $\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{t}^{\text{0}}})$ and $\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{t}^{\text{1}}})$, which are then averaged to obtain the final denoised tomogram. Contrary to those denoising methods, DeepDeWedge fits a network not only to denoise, but also to fill the missing wedge.
@@
Our method is most closely related to IsoNet, which can also do denoising and missing wedge reconstruction. IsoNet takes a small set of, say, one to ten tomograms and produces denoised, missing-wedge corrected versions of those tomograms. Similar as for DeepDeWedge, a randomly initialized network is fitted for denoising and missing wedge reconstruction on sub-tomograms of these tomograms, however the fitting process is different: Inspired by Noisier2Noise, the model is fitted on the task of mapping sub-tomograms that are further corrupted with an additional missing wedge and additional noise onto their non-corrupted versions. After each iteration, the intermediate model is used to predict the content of the original missing wedges of all sub-tomograms. The predicted missing wedge content is inserted into all sub-tomograms, which serve as input to the next iteration of the algorithm.
@@
Different to IsoNet, our denoising approach is Noise2Noise-like, as in CryoCARE. This leads to better denoising performance, as we will see later, as well as requiring fewer assumptions and no hyperparameter tuning. Specifically, Noisier2Noise-like denoising requires knowledge of the noise model and strength (see Supplementary Information 1.2). As this knowledge is typically unavailable, Liu et al. [4] propose approximate noise models from which the user has to choose.
After model fitting, the user must manually decide which iteration and noise level gave the best reconstruction. Thus, IsoNet’s Noisier2Noise-inspired denoising approach requires several hyperparameters for which good values exist but are unknown. Therefore, IsoNet requires tuning to achieve good results. Our denoising approach introduces no additional hyperparameters and does not require knowledge of the noise model and strength.
@@
The main commonality between DeepDeWedge and IsoNet is the Noisier2Noise-like mechanism for missing wedge reconstruction, which consists of artificially removing another wedge from the sub-tomograms and fitting the model to reconstruct the wedge.
Moreover, like IsoNet, DeepDeWedge fills in the missing wedges of the model inputs. In IsoNet, one also has to fill in the missing wedges of the model targets. This is necessary because, contrary to our loss $\ell$ form Equation (2), IsoNet’s loss function does not ignore the targets’ missing wedges via masking in the Fourier domain.
Relation to Other Work.
Another line of works related to DeepDeWedge considers domain-specific tomographic reconstruction methods that incorporate prior knowledge of biological samples into the reconstruction process to compensate for missing wedge artifacts, for example, ICON [18], and MBIR [19]. For an overview of such reconstruction methods, we refer to the introductory sections of works by Ding et al. [20] and Bohning et al. [21]. Liu et al. [4] found that IsoNet outperforms both ICON and MBIR.
DeepDeWedge is also conceptually related to un-trained neural networks, which reconstruct an image or volume based on fitting a neural network to data [22, 23]. Un-trained networks also only rely on fitting a neural network to given measurements. However, they rely on the bias of convolutional neural networks towards natural images [24, 25], whereas in our setup, we fit a network on measurement data to be able to reconstruct from the same measurements.
For cryo-EM-related problems other than tomographic reconstruction, deep learning approaches for missing data reconstruction and denoising have also recently been proposed. Zhang et al. [26] proposed a method to restore the state of individual particles inside tomograms, and Liu et al. [27] proposed a variant of IsoNet to resolve the preferred orientation problem in single-particle cryo-EM [27].
2.3 Theoretical Motivation for the Loss Function
@@
Figure 2: Illustration of a single ground-truth structure $\mathbf{v}^{*}$, and three model input-target pairs corresponding to three different random positions of the original (blue) and additional (red) missing wedge for the setup of our theoretical motivation. For simplicity, we visualize a 2D structure with no additive noise and random in-plane rotations. The inset boxes in each patch show the absolute values of the Fourier transforms of the images and regions that are zeroed out by the missing wedge masks.
@@
Here, we present a theoretical result that motivates the choice of our per-sample loss $\ell$ defined in Equation (2). The discussion in this section does not consider the heuristic of updating the missing wedges of the model inputs, which is part of DeepDeWedge’s model fitting step. Moreover, we consider an idealized setup that deviates from practice to motivate our loss.
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}}=\mathbf{F}^{-1}\mathbf{M}\mathbf{F}\big{(}\mathbf{v}^{*}+\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{n}^{\text{0}}}}\big{)},\qquad{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{1}}}=\mathbf{F}^{-1}\mathbf{M}\mathbf{F}\big{(}\mathbf{v}^{*}+\mathbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{n}^{\text{1}}}}\big{)},$ (4)
@@
where $\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{n}^{\text{0}}}},\mathbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{n}^{\text{1}}}}\in\mathbb{R}^{N\times N\times N}$ are random noise terms and $\mathbf{M}\in\{0,1\}^{N\times N\times N}$ is the missing wedge mask. From the first measurement, we generate a noisier observation ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}}=\mathbf{F}^{-1}\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}\mathbf{F}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}}$ by applying a second missing wedge mask $\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}$. The noisier observation has two missing wedges: the wedge introduced by the first and the wedge introduced by the second mask. We assume that the two masks follow a joint and symmetric distribution, e.g., that for each mask, a random wedge is chosen uniformly at random. Figure 2 illustrates three data points in 2D.
@@
We then train a neural network $\mathrm{f}_{\text{$\bm{\theta}$}}$ to minimize the loss
@@
$\mathrm{L}(\text{$\bm{\theta}$})=\mathop{\mathbb{E}}_{\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{n}^{\text{0}}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{n}^{\text{1}}}}\left[\left|\left|\big{(}\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}\mathbf{M}+2\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}^{C}\mathbf{M}\big{)}\mathbf{F}\Big{(}\mathrm{f}_{\text{$\bm{\theta}$}}\big{(}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}}\big{)}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{1}}}\Big{)}\right|\right|_{2}^{2}\right],$ (5)
@@
After training, we can use the network to estimate the ground-truth volume by applying the network to another noisy observation ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}}$. The following proposition, whose proof can be found in Supplementary Information 3, establishes that this is equivalent to training the network on a supervised loss to reconstruct the input ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}}$, provided the two masks are non-overlapping.
@@
Assume that the noise $\mathbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{n}^{\text{1}}}}$ is zero-mean and independent of the noise $\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{n}^{\text{0}}}}$, and of the masks $(\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}})$, and assume that the noise $\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{n}^{\text{0}}}}$ is also independent of the masks. Moreover, assume that the joint probability distribution $P$ of the missing wedge masks $\mathbf{M}$ and $\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}$ is symmetric, i.e., $\mathrm{P}(\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}})=\mathrm{P}(\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}},\mathbf{M})$, and that the missing wedges do not overlap. Then the loss $\mathrm{L}$ is proportional to the supervised loss
$\mathrm{R}(\text{$\bm{\theta}$})=\mathop{\mathbb{E}}_{\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{n}^{\text{0}}}}\left[\left|\left|\mathrm{f}_{\text{$\bm{\theta}$}}\big{(}\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}}}\big{)}-\mathbf{\mathbf{v}_{missing}^{*}}\right|\right|_{2}^{2}\right],$ (6)
i.e., $\mathrm{L}(\text{$\bm{\theta}$})=\mathrm{R}(\text{$\bm{\theta}$})+c,$ where $c$ is a numerical constant independent of the network parameters $\bm{\theta}$.